the sum of three numbers is 99.the first is 10 greater than second and 20 greater than the third.find the largest of the three numbers?
Answer by oberobic(2304)

There are 3 numbers: x, y, z.
x+y+z=99
y = x-10
z = x-20
.
x + x-10 +x-20 = 99
3x -30 = 99
3x = 129
x = 43
y = 33
z = 23
.
x + y + z = 43 +33 + 23
43 + 33 + 23 = 99
Check.
.
The greatest of the three numbers is 43.
.
Done.